# TAP4LLM: Table Provider on Sampling, Augmenting, and Packing Semi-structured Data for Large Language Model Reasoning

### 14 Dec 2023 v1

@@ 3.1.1. Rule-based Sampling
3.1.2. Embedding-based Sampling
Embedding-based sampling is an advanced approach to table sampling based on high-dimensional vector embeddings. Instead of adhering to strict rules or criteria in rule-based sampling, embedding-based methods leverage the semantic and contextual representation of each row and column. By mapping each row or column to vectors, this method harnesses the power of spatial relationships within the embedding space to guide sampling decisions. Specifically, we propose two embedding-based sampling methods as follows:
(1) Query-based Sampling: Query-based Sampling is a tailored approach emphasizing the semantics relevance of table cells to the utterance. The process is exactly illustrated in Eq. 1. Note that the default query-based sampling is the row-based method. We also study the significance of column grounding in table sampling by testing query-based sampling in conjunction with column grounding, as shown in Table 3.1. By incorporating column grounding, the sampling process revolves around both rows and columns, offering a dynamic table subset where both rows and columns are susceptible to change based on the user query.
(2) Clustering-based Sampling: In Clustering-based Sampling, the core principle is to group similar rows or columns based on their embedding representation. Specifically, let $T$ be a table where $R_{T}$ is the set of rows and $C_{T}$ is the set of columns. Let $E:R_{T}\cup C_{T}\rightarrow\mathbb{R}^{d}$ be an embedding function that maps each row or column to a $d$-dimensional vector by capturing its semantic content. The goal of clustering-based sampling is to extract a sub-table $T^{\prime}\subseteq T$ and ensure the preservation of data diversity. We use K-Means (MacQueen et al., 1967) as our clustering-method backbone. Specifically, for each $r\in R_{T}$ or $c\in C_{T}$, compute its embedding as $e=E(r)$ or $e=E(c)$. Then, utilize the K-Means algorithm, and partition the set of embeddings into $n$ clusters: $\{C_{1},C_{2},\ldots,C_{n}\}$. For each cluster, $C_{i}$, select the top-$K$ rows or columns based on a specified criterion (e.g., closeness to centroid). At last, aggregate the chosen rows or columns from all clusters to form $T^{\prime}$.
3.2. Table Augmentation
Table augmentation has emerged as a crucial approach in navigating the challenges associated with the intricate and varied nature of tabular data. Two primary benefits arise from this approach. Firstly, Enhanced Contextual Understanding: Raw tables often come without ample context or explicit metadata, such as distribution or dimension/measure categorizations for columns. By supplementing tables with metadata and attributes, we can achieve a more profound grasp of the table’s intrinsic structure and semantics and further enrich the tabular data. Secondly, Bridging External Knowledge Gaps: tables alone might not encompass all the required information to provide comprehensive answers to certain queries. By retrieving external knowledge from reliable sources, e.g., Wikipedia, we can aid the language models in understanding the broader context of the query, leading to more informed and nuanced responses. (See the various knowledge aspects used in TAP4LLM from Table 1.)
3.2.1. Metadata-based Augmentation
Metadata is defined as a form of background knowledge to understand the field semantics for correctly operating on table fields (or columns) and to further find common patterns in daily analysis (He et al., 2023). This analytical knowledge, particularly of field semantics, is able to increase the applicability across various tasks. Specifically, we consider the following metadata-based augmentation types:
(1) Dimension / Measure: This is one type of metadata used in Tableau (Hoelscher and Mortimer, 2018) and Excel (Ding et al., 2019) across diverse features. As the name suggests, the method involves categorizing each field in a table as either measure or dimension. The measure contains numerical data that can be subjected to calculations, such as the “Price” and “Discount”. The dimension provides categorical information used for filtering, grouping, and labeling, such as the “Product Name” and “Category”. Correctly classifying fields as either a measure or a dimension is crucial to determining feasible operations on the data and influences the accuracy and relevance of data analysis.
(2) Semantic Field Type: Besides identifying whether a field is a measure or a dimension, semantic field type specifies the meaning and format of the data within each field based on knowledge graphs. For example, the dimension field includes semantic field types such as “Consumer Product” and “Category”, etc. Measure field includes semantic field types such as “Money” and “Ratio”, etc. We follow the work  (He et al., 2023) as a reference to this term.
(3) Table Size: The size of a table is defined by its number of rows and columns. It provides essential context when determining the computational complexity of operations or understanding data density and granularity.
(4) Statistics Feature: Statistics feature provides a quantitative representation of the tabular data. These features serve as numerical descriptors that summarize key aspects of the table datasets, aiding LLMs in understanding the overall characteristics and tendencies. Generally, statistics features include four categories (He et al., 2023): (a) Progression features (b) String features (c) Number range features (d) Distribution features, discussed in Section §6. We conducted empirical studies on common statistical features to identify the most appropriate combination for optimal utilization of  TAP4LLM.
(5) Header Hierarchy: Tables are often used to present data in a structured format, and headers play a crucial role in defining the meaning and context of the data in each column or row. The header hierarchy typically includes different levels of headers, each providing a level of organization and categorization for the data.
3.2.2. Retrieval-based Augmentation
Large Language Models have occasionally been observed to generate hallucinated or factually inaccurate text (Zhou et al., 2021; Zhao et al., 2023). To mitigate these issues, several works have proposed to strengthen LLMs with information retrieval systems (retrieval-augmented LLMs) (Shi et al., 2023; Jiang et al., 2023; Nakano et al., 2022), which enables LLMs to retrieve relevant contents from an external repository (knowledge corpus). It has been verified that retrieval-augmented LLMs can generate texts in response to user input with fewer hallucinations (Nakano et al., 2022). Furthermore, by incorporating customized private data resources, retrieval-augmented LLMs can respond to in-domain queries that cannot be answered by LLMs trained with public data. As previous works (Nakano et al., 2022; Shi et al., 2023; Jiang et al., 2023) suggest, LLMs can generate more factual answers by feeding the references retrieved from the external corpus. In TAP4LLM, we have fortified the document retrieval capabilities of LLMs and consider two components: (i) document references: to provide supplemental relevant web pages as the references for the given table; (ii) term explanation: to explain strange/ambiguous term in the given table. We utilize technologies like vector databases (Wang et al., 2021b) and LangChain (LangChain, 2022) to facilitate the retrieval of pertinent information from Wikipedia [^3] https://www.wikipedia.org/.
Table 2. LLM-based Cell Selection Criteria and Exact Prompt Template.
Criteria
Description
Cell Position
Cell Content
Cell Formatting
Cell Context
Cell Properties
(1) Docs References: This process involves associating tables with relevant documents or sources for in-depth insights or references. For example, suppose we have a table titled “2023 Fortune 500 Companies”. This table contains various information about the top 500 companies as ranked by Fortune in 2023, including their revenue, number of employees, and market capitalization. Docs references could fetch the actual 2023 Fortune 500 list from the Fortune website, Wikipedia pages discussing the Fortune 500 concept and its criteria, or analytical articles discussing the companies on the 2023 list. In our setting, we leverage Langchain (LangChain, 2022) to retrieve wiki pages from wikipedia.org. We craft queries by concatenating the table header and the table’s title into a single string. These queries are then used to identify and fetch the relevant Wikipedia pages, which act as informative document references in our study.
(2) Term Explanation: Compared to the docs references, term explanation focuses on providing definitions and explanations for specific strange terms or values in the table cells. For example, if a cell mentions a technical term or an acronym, the term explanation module could source a brief definition or background from reliable web sources (such as Wikipedia, wolfram,etc) on that term, ensuring that the strange term will not be forwarded to LLMs. To ensure the efficacy and accuracy of term explanations, we introduce two distinct approaches for selecting the cell that is required to be explained: i) LLM-based Cell Selection Module: To pinpoint the exact cell warranting explanation, we harness the capabilities of LLMs. The selection prompt is meticulously constructed, taking into account various factors including: (1) Cell Position; (2) Cell Content; (3) Cell Formatting; (4) Cell Context; (5) Cell Properties. A detailed description and the specific prompt utilized to determine which cells require explanation can be found in Table 2. ii) Heuristics-based Cell Selection: Inspired by the methodology presented in (Herzig et al., 2020), we introduce a heuristics-based cell selection, which is predicated upon the following criteria: (1) Explicit Mention: whether the cell’s value is explicitly referenced in the query. (2) Comparative Value: whether the cell’s value is greater or less than a value mentioned in the query. (3) Superlative Value: whether the cell’s value represents a maximum or minimum across the entire column, especially when the query incorporates superlative terms. The comparative experiment results of these two variants can be found in Table 6.
3.2.3. Self-consistency-based Augmentation
We follow (Sui et al., 2023) to provide the self-consistency-based augmentation approach. First, we append the instruction “Identify critical values and ranges of the last table related to the statement” to the initial prompt, and then forward this prompt to the LLM. The output generated from this instruction is then incorporated back into the prompt. Following this, we reintroduce the enriched prompt, now containing both the initial query and the newly generated insights, to the LLM along with the task-specific instructions for further processing. This iterative approach, namely self Prompting, allows for a deeper contextual understanding and analysis by LLMs. Formally, self-prompting is a simple idea that utilizes prompting twice to leverage the capabilities of LLMs in understanding structured data. In essence, through this self-consistency-based augmentation, we seek to harness and amplify the latent structured data reasoning prowess of LLMs.
3.3. Table Packing
Table 3. The distribution of the used datasets.
Property SQA FEVEROUS TabFact HybridQA ToTTo Spider
Unique Query (Set Size) 1,228 1,322 9,228 6,268 8,026 10,181
Unique Table 432 942 1,342 4,364 5,934 500
SQL Query - - - - - 5,693
Columns per tables (Median/Avg) 4 / 6.4 4 / 5.5 4 / 5.5 4 / 4.3 6 / 8.8 4 / 4.5
Domain Wikipedia Wikipedia Wikipedia Wikipedia Wikipedia Wikipedia
Evaluation Metric Exact Match Exact Match Exact Match Exact Match BLEU-4 Exact Match
[…]
Moreover, our empirical study (Sui et al., 2023) emphasizes a noteworthy observation regarding using markup languages like HTML or XML leads to much better generation quality by LLMs over TQA and TFV. In this pattern, TAP4LLM support multiple serialization functions, e.g., HTML, XML, JSON, CSV, NL+Sep (one of the typical options, e.g., using ‘||||’ as cell/column separator) and Markdown,etc.
4. Experiments
4.1. Experiment Settings
4.1.1. Downstream Tasks and Datasets
@@
In this paper, we mainly focus on tabular reasoning with two major tasks: TQA & TFV. We conduct experiments on five typical datasets and the distribution of the datasets can be found in Table 3. In addition, to extend our work to databases containing table structures, we also set up  TAP4LLM on Spider  (Yu et al., 2018) dataset. Specifically, we use: (1) SQA, which is constructed by decomposing a subset of a highly compositional dataset, WTQ (Pasupat and Liang, 2015). The dataset consists of 1,288 unique queries corresponding to 432 tables, with each table having 18.5 rows and 6.4 columns on average; (2) HybridQA, which is designed as a large-scale multi-hop question-answering dataset over heterogeneous information of both structured tabular and unstructured textual forms. The dataset consists of 6,268 unique questions and each question is aligned with a Wikipedia table. Compared to the SQA dataset, HybridQA has shorter column numbers, which facilitates the understanding of the table’s structure boundaries. (3) ToTTo is a high-quality English table-to-text dataset. It proposes a controlled generation task that involves synthesizing a one-sentence description given a Wikipedia table and a set of highlighted table cells. The dataset contains 8,026 samples, each comprising a Wikipedia table with highlighted cells. Each table contains 16 rows and 6 columns on average. (4) FEVEROUS is a fact verification dataset over structured information. The dataset consists of 1,322 verified claims. Each claim is annotated with evidence in the form of sentences and cells from tables in Wikipedia. Each annotation also includes a label indicating whether the evidence supports, refutes, or does not provide enough information to make a decision. Each table contains 26.3 rows and 5.5 columns on average. (5) TabFact is another fact verification dataset where the tables are extracted from Wikipedia and the sentences are composed by crowd workers. Compared to the FEVEROUS dataset, TabFact encompasses a larger number of samples and each table has fewer rows, has 14 rows per table on average. (6) Spider is a semantic parsing text-to-SQL dataset across multiple domains. The dataset consists of 10,181 questions and 5,693 unique complex SQL queries on 200 databases with multiple tables.
4.1.2. Models
@@
In this study, we evaluate the performance of the recent dominant LLM models, 1) Instruct-GPT-3.5 (Ouyang et al., 2022), using versions gpt-3.5-turbo, gpt-3.5-turbo-16k; 2) GPT-4, using the latest version of gpt-4 model. Unless otherwise specified, we utilize gpt-3.5-turbo in all experiments. In the sampling methods, we use text-embedding-ada-002 (ope, [n.d.]) for row and query embedding generation. The comparison experiments using other embeddings models, such as, text-search-ada-doc-001, bge-largen-en (Xiao et al., 2023), all-MinLM-L6-v2 (Reimers and Gurevych, 2019) can be found in Table 7. The development of TAP4LLM begins with the foundation provided by LLMs. In designing our framework, we opt to use OpenAI models as our base model due to their excellent capabilities in language reasoning. However, the choice is not exclusive. Since TAP4LLM use natural language as an intermediary for interactive communication between the table and LLMs, it can also support other outstanding open-sourced models using natural language as input, such as LLaMa (Touvron et al., 2023), Phoenix (Chen et al., 2023), ChatGLM (Zeng et al., 2022), Ziya (IDEA-CCNL, 2023), and Baichuan (Intelligence, 2023). This design provides versatility and flexibility in TAP4LLM implementation.
Table 4. Comparative results of the table sampling methods. The term “w/ Column Grounding” refers to the method consider both row-based and column-based sampling (sometimes referred to as “grounding”). “GPT-3.5” refers to the OpenAI released model gpt-3.5-turbo-32k, with 32k token-sized context window; In contrast, “GPT-3.5 truncated” refers to gpt-3.5-turbo, with 4k token-sized context window, where most tables will be truncated according to this token limitation. The top-3 performances on each dataset are highlighted in green, with the best performance being both bold and underlined.
Table Samping Methods SQA FEVEROUS TabFact HybridQA ToTTo Spider
Clustering-based Sampling 28.10% 63.50% 55.40% 24.03% 48.30% 77.43%
Query-based Sampling 28.32% 63.32% 59.80% 24.32% 49.14% 80.27%
  w/ Column Grounding 29.12% 64.74% 60.23% 25.14% 53.42% 81.03%
Random Sampling 27.30% 60.30% 55.17% 23.60% 40.12% 74.58%
Evenly Sampling 26.72% 61.87% 54.63% 5.32% 29.41% 72.03%
Content Snapshot 28.24% 63.10% 56.92% 23.40% 47.51% 78.93%
No sampling (GPT-3.5) 27.60% 60.12% 56.20% 14.10% 47.42% 72.15%
No sampling (GPT-3.5, truncated) 23.54% 43.54% 52.12% 23.12% 30.42% 68.47%
Table 5. Comparative results of the table augmentation methods. We use query-based sampling method (no augmentation) as the baseline. The term “Delta” refers to the performance gap between each method and the baseline. The top-3 performance gap on each dataset are highlighted in green, with the best performance being underlined. Noted that since only the Totto dataset contains hierarchical headers, we only provide the “header hierarchy” method on this dataset.
Table Augmentation Methods SQA FEVEROUS TabFact HybridQA ToTTo Spider
Acc Delta Acc Delta Acc Delta Acc Delta BLEU-4 Delta Acc Delta
baseline 28.32% 0.00% 63.32% 0.00% 59.80% 0.00% 24.32% 0.00% 49.14% 0.00% 80.27% 0.00%
Dimension/Measure+ Semantic Field Type 30.12% 1.80% 65.72% 2.40% 62.67% 2.87% 26.12% 1.80% 51.25% 2.11% 82.45% 2.18%
Table Size 28.85% 0.53% 63.40% 0.08% 60.30% 0.50% 24.94% 0.62% 49.03% -0.11% 79.86% -0.41%
Statistics Feature 31.22% 2.90% 66.51% 3.19% 62.33% 2.53% 26.13% 1.81% 50.57% 1.43% 80.94% 0.67%
Header Hierarchy - - - - - - - - 48.64% -0.50% - -
Docs References 33.45% 5.13% 63.13% -0.19% 61.32% 1.52% 25.12% 0.80% 52.74% 3.60% 77.65% -2.62%
Term Explanations 31.59% 3.27% 64.12% 0.80% 62.32% 2.52% 26.24% 1.92% 53.21% 4.07% 80.48% 0.21%
Self Prompting 30.45% 2.13% 65.24% 1.92% 62.32% 2.52% 26.64% 2.32% 52.36% 3.22% - -
4.2. Comparison Results of Table Sampling
According to Table 4, we conduct the comparative experiments on multiple table sampling methods and make several observations as follows:
(1) Effectiveness of query-based and clustering-based Sampling: The query-based sampling method, combined with column grounding, achieves the best performance over all datasets. It indicates that sampling specific, relevant parts of the table that are relevant to the query can enhance LLMs’ capability to produce more accurate results. Additionally, the use of both row-based and column-based sampling can provide a fine-grained subset of the table that covers the most essential information for the reasoning process. Similarly, the clustering-based sampling method demonstrates satisfactory performance. This method uses the patterns and similarities within the data point in each table to create the clusters that guide the sampling process. It doesn’t consider the relevance between the table and the query. Instead, it mainly focuses on the inner information of the table across different clusters.
(2) Potential of Content Snapshot: “Content Snapshot (Yin et al., 2020)” shows result that is relatively close to “Clustering-based Sampling” and “Query-based Sampling”. It indicates that capturing a snapshot of essential information from the table can provide a useful cross-section for the LLMs to learn from. Even though it might not always capture the most relevant or critical information that is needed for table reasoning, it demonstrates its potential as a viable sampling technique. Furthermore, compared to the query-based or clustering-based sampling, this method doesn’t require time-consuming embedding generation. Instead, it calculates the $n$-gram overlap with the utterance and the process is linear with the number of rows.
(3) Table Sampling vs. Direct Encoding (Truncation): Directly leveraging the GPT-3.5-turbo model with 32k tokens limitation shows worse performance compared to using table sampling methods. Even though it can cover more information on the table, however, the results demonstrate that encoding much of the table information might introduce noise and disrupt the table reasoning process. Furthermore, using a 4K token-sized context window model with truncation results in even poorer performance compared to the model with a 32k token limitation. It indicates that truncating the table could lead to the loss of vital context or information, potentially restricting the model’s capability to fully understand the table and generate comprehensive answers.
4.3. Comparison Results of Table Augmentation.
For the comparative experiments of table augmentation methods, we use the query-based sampling method as the baseline method and report the performance gap between each augmentation method and the baseline. According to Table 5, several insights can be found as follows:
Table 6. Comparative results of cell selection functions. We use query-based sampling as the baseline.
Methods SQA FEVEROUS TabFact HybridQA ToTTo Spider
Acc Delta Acc Delta Acc Delta Acc Delta BLEU-4 Delta Acc Delta
LLM-based 31.59% 3.27% 64.12% 0.80% 62.32% 2.52% 26.24% 1.92% 53.21% 4.07% 80.48% 0.21%
Heuristics-based 29.59% 1.27% 63.72% 0.40% 61.58% 1.78% 25.24% 0.92% 51.21% 2.07% 80.33% 0.06%
(1) Effectiveness of table augmentation: From the experimental results, we found that table augmentation methods further improve LLM’s reasoning ability after sampling. For example, ”Dimension/Measure + Semantic Field Type” achieves higher accuracy across all six datasets (most significant increase on TabFact +2.87%). ”Statistics Feature” performs effectively on datasets that include a higher proportion of statistical cell contents (FEVEROUS +3.19% ). ”Docs References” and ”Term Explanations” add meaningful context and semantic understanding to the model’s processing of tables, with (SQA +5.13%, ToTTo +4.07% ). Also, ”self-prompting” helps the model generate better queries or responses by encouraging it to use its own outputs as a guide for further response generation (ToTTo +3.22%). We also found some methods might not improve the model’s performance by adding extra information. According to the result, the model performance with ”Table Size” is minimal, suggesting little help in adding substantial semantic information. In addition, the result of ”Header Hierarchy” shows that introducing a hierarchy may complicate the model’s ability to process the tabular information in some contexts, possibly by adding unnecessary complexity.
(2) Comparison of different cell selection methods: For the ”Term Explanations” method, we compare two distinct cell selection functions as shown in Table 6. We find that LLM-based cell selection outperforms the heuristics-based cell selection with improvements in “Delta” ranging from 0.80% to 4.07%. While achieving higher performance, the LLM-based method also increases the calling budget as it requires additional LLM calls. These results indicate that the method’s effectiveness varies with the dataset. i.e. It’s beneficial for datasets requiring complex text understanding and generation (SQA and ToTTo). However, its impact is less distinct or even slightly negative in datasets involving different types of data or nuanced tasks (FEVEROUS and HybridQA).
(3) Combination of augmentation methods: Through the experiment results, we have investigated the best combination of single sampling method and augmentation method on each dataset. We also observe that different methods perform well on the same dataset. For example, ”Dimension/Measure + Field Type”, ”Statistics Feature”, ”Term Explanation” and ”Self Prompting” all show significant improvement on the TabFact dataset. This suggests additional research on the superposable effects through combining multiple augmentation methods. We leave this work for future investigation.
4.4. Comparison Results of Embedding Type.
@@ 4.2. Comparison Results of Table Sampling
Table 7. Comparative results of different embedding models on query-based sampling method without any augmentation method. We use all-MinLM-L6-v2 for the sentence-transformer. The highest performance of each dataset is bold.
@@
bge-large-en (Xiao et al., 2023) 26.76% 62.87% 56.31% 22.65% 47.32% 78.25%
sentence-transformer (Reimers and Gurevych, 2019) 26.32% 63.31% 58.94% 23.78% 50.12% 80.05%
@@
Table 8. Comparative results of various types of statistical features. The experiment setting is the same as Section 5. The highest performance of each dataset is bold.
Statistics Features Type SQA FEVEROUS TabFact HybridQA ToTTo Spider
Progression features 29.20% 64.26% 60.45% 25.11% 49.53% 77.47%
String features 28.56% 63.13% 61.38% 24.83% 48.29% 73.56%
Number range features 29.13% 62.18% 59.03% 24.53% 49.68% 76.32%
Distribution features 30.28% 66.34% 62.18% 24.76% 49.34% 79.14%
Statistics features 31.22% 66.51% 62.33% 26.13% 50.57% 80.94%
4.5. Comparison Results of Statistics Features

### 17 Feb 2024 v2

@@ D.1 Downstream Tasks and Datasets
In this paper, we mainly focus on tabular reasoning with two major tasks: TQA & TFV. We conduct experiments on five typical datasets and the distribution of the datasets can be found in Table 5. In addition, to extend our work to databases containing table structures, we also set up  TAP4LLM on Spider  Yu et al. (2018) dataset. Specifically, we use: (1) SQA Iyyer et al. (2017), which is constructed by decomposing a subset of a highly compositional dataset, WTQ Pasupat and Liang (2015). The dataset consists of 1,288 unique queries corresponding to 432 tables, with each table having 18.5 rows and 6.4 columns on average; (2) HybridQA Chen et al. (2020b), which is designed as a large-scale multi-hop question-answering dataset over heterogeneous information of both structured tabular and unstructured textual forms. The dataset consists of 6,268 unique questions and each question is aligned with a Wikipedia table. Compared to the SQA dataset, HybridQA has shorter column numbers, which facilitates the understanding of the table’s structure boundaries. (3) ToTTo Parikh et al. (2020) is a high-quality English table-to-text dataset. It proposes a controlled generation task that involves synthesizing a one-sentence description given a Wikipedia table and a set of highlighted table cells. The dataset contains 8,026 samples, each comprising a Wikipedia table with highlighted cells. Each table contains 16 rows and 6 columns on average. (4) FEVEROUS Aly et al. (2021) is a fact verification dataset over structured information. The dataset consists of 1,322 verified claims. Each claim is annotated with evidence in the form of sentences and cells from tables in Wikipedia. Each annotation also includes a label indicating whether the evidence supports, refutes, or does not provide enough information to make a decision. Each table contains 26.3 rows and 5.5 columns on average. (5) TabFact Chen et al. (2020a) is another fact verification dataset where the tables are extracted from Wikipedia and the sentences are composed by crowd workers. Compared to the FEVEROUS dataset, TabFact encompasses a larger number of samples and each table has fewer rows, has 14 rows per table on average.
Metrics. For TQA and TFV tasks (SQA, FEVEROUS, TabFact and HybridQA), we report the exact match accuracy of answer sets. For data-to-text generation task (ToTTo), we report BLEU-4 score.
D.2 Models
In this study, we evaluate the performance of the recent dominant LLM models, 1) Instruct-GPT-3.5 Ouyang et al. (2022), using versions gpt-3.5-turbo, gpt-3.5-turbo-16k; 2) GPT-4, using the latest version of gpt-4 model; 3) Llama-2-70B  Touvron et al. (2023), using version 17; 4) Mixtral-8x7B  Jiang et al. (2024), using version 0.1.
@@ D.2 Models
Unless otherwise specified, we utilize gpt-3.5-turbo in all experiments. In the sampling methods, we use text-embedding-ada-002 ope for row and query embedding generation. The comparison experiments using other embeddings models, such as, text-search-ada-doc-001, bge-largen-en Xiao et al. (2023), all-MinLM-L6-v2 Reimers and Gurevych (2019) can be found in Table 7. The development of TAP4LLM begins with the foundation provided by LLMs. In designing our framework, we opt to use OpenAI models as our base model due to their excellent capabilities in language reasoning. However, the choice is not exclusive. Since TAP4LLM use natural language as an intermediary for interactive communication between the table and LLMs, it can also support other outstanding open-sourced models using natural language as input, such as Phoenix Chen et al. (2023), ChatGLM Zeng et al. (2022), Ziya IDEA-CCNL (2023), and Baichuan Intelligence (2023). This design provides versatility and flexibility in TAP4LLM implementation.
Appendix E Additional Experiments
E.1 Comparison Results of Embedding Type.
@@ Appendix E Additional Experiments
Table 7: Comparative results of different embedding models on query-based sampling method without any augmentation method. We use all-MinLM-L6-v2 for the sentence-transformer. The highest performance of each dataset is bold.
@@
bge-large-en Xiao et al. (2023) 26.76% 62.87% 56.31% 22.65% 47.32% 78.25%
sentence-transformer Reimers and Gurevych (2019) 26.32% 63.31% 58.94% 23.78% 50.12% 80.05%
@@
E.2 Comparison Results of Statistics Features